OpenStudy (anonymous):

Using double and identities, find the exact solutions for 0

OpenStudy (xapproachesinfinity):

let met try, i haven't done this in awhile sin2x=root3sinx 2sinxcosx=root3sinx divide by sinx 2cosx=root3 divide by 2 cosx=root3/2 x=pi/6, 5pi/6

OpenStudy (xapproachesinfinity):

i hope that's the anwser, i got rusty in this

OpenStudy (anonymous):

Thank you so much! I thought sin2x was sinx + cosx, (silly mistake, thats pythagorean) not multiplying the 2.

OpenStudy (xapproachesinfinity):

actually the solution are x=0. pi, pi/6

OpenStudy (xapproachesinfinity):

i just checked it by unit circle

OpenStudy (xapproachesinfinity):

you're welcome

OpenStudy (anonymous):

how did you get x=0 and pi?

OpenStudy (xapproachesinfinity):

Sorry yesterday i went to sleep Here the steps before i missed a step and missed some solutions because it isn't allowed to divide by sinx since it could equal zero too which not allowed in the denumerator so it is as follwed: sin2x=root3sinx substract root3sinx from both sides -root3sinx+sin2x=0 use the indentity sin2x=2sinxcosx -root3sinx+2sinxcosx=0 factor sinx sinx(-root3+2cosx)=0 then sinx=0 or -root3+2cosx=0 sinx=0 or 2cosx=root3 sinx=0 or cosx=root3/2 now solve we get x=0, pi or x=pi/6 we cannot go further since the interval is 0<= x <= pi